Twice the sum of a number and 3 is the same as 1 subtracted from the number. Find the number

Answer by math_helper(2461)

Let x = the number

2(x+3) is "twice the sum of a number and three"
x-1 is "1 subtracted from the number"
Expanding 2(x+3) and setting the two equal to each other:
2x+6 = x-1
x = -7


Ans: x = -7

Check:
2(-7+3) = 2(-4) = -8
-7-1 = -8 (ok)